### v1

@@ 3.1 Main Theorem
Theorem 3.1
Suppose Sections 3 and 3 hold, and $\mathcal{Z}^{*}(X)$ defined in Section 3 is a singleton almost surely. Suppose there exists a positive number $\tilde{r}$ (that depends on $n$) that upper bounds the critical radii of the function classes $\mathcal{G}_{1},\mathcal{G}_{2}$ almost surely (i.e., $\mathcal{R}_{n/2}(\mathcal{G}_{1},\tilde{r})\leq{\tilde{r}}^{2}$ and $\mathcal{R}_{n/2}(\mathcal{G}_{2},\tilde{r})\leq{\tilde{r}}^{2}$) and satisfies the inequalities $3n\tilde{r}^{2}/128\geq\log\log_{2}(1/\tilde{r})$, and $42\exp\left(-3n\tilde{r}^{2}/128\right)\leq\delta/$ 4. Then, there exists a positive constant $\tilde{C}(\alpha,\gamma)$ such that with probability at least $1-\delta$, we have
@@
The upper bound in Theorem 3.1 involves several different types of error terms. The first type is the critical radii $\tilde{r}$ that characterize the complexity of the function classes $\mathcal{G}_{1},\mathcal{G}_{2}$. This is a common complexity measure for function classes in statistics and machine learning Wainwright (2019). For example, as we will discuss below, the critical radii scale as $\tilde{O}(\sqrt{\eta/n})$ if $\mathcal{G}_{1},\mathcal{G}_{2}$ are VC subgraph classes of dimension $\eta$. The second type is the term $\operatorname{Rate}^{\textbf{N}}(n/2,\delta/4)$ resulted from the errors in estimating the nuisance functions $f_{0},\Sigma_{0}$. Similar nuisance estimation errors also appear in the previous literature on offline contextual bandit learning Athey and Wager (2021), Zhou et al. (2023), Chernozhukov et al. (2018) or more general learning problems that involve nuisance functions Foster and Syrgkanis (2023). Below we will further discuss this term for different choices of score functions $\theta$. The third type is the misspecification error term $\text{Reg}\left(\tilde{\pi}^{*}\right)$. It is natural to expect a bigger decision regret when the misspecification error $\text{Reg}\left(\tilde{\pi}^{*}\right)$ is higher. In particular, when the policy class $\Pi_{\mathcal{F}}$ is well-specified, i.e., $\pi_{f_{0}}\in\Pi_{\mathcal{F}}$, we would have $\text{Reg}\left(\tilde{\pi}^{*}\right)=0$, and the regret upper bound would scale with the function class complexity through a fast rate $O(\tilde{r}^{\frac{2\alpha+2}{\alpha+2}})$. For VC subgraph classes with $\tilde{r}=\tilde{O}(\sqrt{\eta/n})$, the rate can range from $\tilde{O}(\sqrt{\eta/n})$ for $\alpha=0$ to $\tilde{O}({\eta/n})$ for $\alpha=\infty$. However, if the policy class is misspecified so that $\text{Reg}\left(\tilde{\pi}^{*}\right)>0$, then the dominating term related to $\tilde{r}$ would be a slow rate $O(\tilde{r})$. This reveals an interesting phase transition between the correctly specified and misspecified settings, which is not discovered in the previous theory that considers only well-specification Hu et al. (2022).
Remark 3.2
The constant coefficients in the regret upper bound can be improved when $2(1+\alpha)/\alpha$ is an integer (which accommodates the case of $\alpha=1$ justified in Hu et al. (2024)):
@@
Notably, the constant in front of the misspecification error $\text{Reg}\left(\tilde{\pi}^{*}\right)$ becomes $1$ instead of $2$, which we believe is tight. This upper bound follows from nearly the same proof as Theorem 3.1 except that it handles an inequality slightly differently. Specifically, the proof of Theorem 3.1 involves a transcendental inequality of the form $\text{Reg}(\hat{\pi})\leq c_{1}\text{Reg}(\hat{\pi})^{\frac{\alpha}{2(1+\alpha)}}\tilde{r}+c_{2}$ for certain positive terms $c_{1},c_{2}$. This inequality is difficult to solve exactly, so we can only get an upper bound on its solution. It turns out that we can get a better upper bound when $2(1+\alpha)/\alpha$ is an integer.
@@
Proposition 3.3
Suppose that the mean square errors of the nuisance estimates satisfy that with probability $1-\delta$ over the randomness of the nuisance sample:
@@
Compared to the DM and ISW scores, the impact of the estimation errors of the nuisances in the DR score is of only second order, i.e., $O(\chi_{n,\delta}^{2})$ instead of $O(\chi_{n,\delta})$. This echos the benefit of DR methods in causal inference and offline contextual bandit learning Athey and Wager (2021), Chernozhukov et al. (2018, 2019). Notably, here we only need to bound the projected error on the nuisance estimator $\hat{f}$, which handles the setting when $\text{span}(\mathcal{Z})$ does not cover the whole $\mathbb{R}^{d}$ space, as is the case with the shortest path problem in Section 5.
@@ 3.2 Byproduct: Fast Rates in the Full Feedback Setting with Misspecification
Theorem 3.5
Suppose $\mathcal{Z}^{*}(X)$ defined in Section 3 is a singleton almost surely. Define
@@
and $\tilde{r}^{\text{F}}$ be any solution to the inequality $\mathcal{R}_{n}(\mathcal{G}^{\text{F}},r)\leq r^{2}$ satisfying $3n(\tilde{r}^{\text{F}})^{2}/64\geq\log\log_{2}(1/\tilde{r}^{\text{F}})$ and $2\exp\left(-3n(\tilde{r}^{\text{F}})^{2}/64\right)\leq\delta$. If Section 3 further holds, then, with probability at least $1-\delta$,
@@
The regret bound is similar to that in Theorem 3.1 for the bandit setting, except that it does not have the nuisance error term $2\text{Rate}^{\textbf{N}}(n/2,\delta/4)$. This is because in the full feedback setting we observe the whole $Y$ vector, so we do not need to estimate any nuisance. When $\Pi_{\mathcal{F}}$ is well-specified, i.e., $\text{Reg}\left(\tilde{\pi}^{*}\right)=0$, and $\tilde{r}^{\text{F}}=\tilde{O}(1/\sqrt{n})$, the bound in Theorem 3.5 recovers the fast rate $\tilde{O}(n^{-\frac{\alpha+1}{\alpha+2}})$ in Hu et al. (2022).
@@ 4 Computationally Tractable Surrogate Loss
The IERM objective is generally nonconvex in $f\in\mathcal{F}$, making it computationally intractable to optimize. In the full feedback setting, tractable surrogate losses have been proposed (Elmachtoub and Grigas 2022, Huang and Gupta 2024). In this section, we briefly explain the SPO+ loss in Elmachtoub and Grigas (2022) and how it can be used in the bandit setting.
@@
This IERM loss is equivalent to the “smart predict-then-optimize” (SPO) loss in Definition 1 of Elmachtoub and Grigas (2022). Letting $z^{*}(y)\in\arg\min_{z\in\mathcal{Z}}y^{\top}z$ with the same tie-breaking rule as in $\pi_{f}$, Elmachtoub and Grigas (2022) propose the SPO+ surrogate loss:
@@ 5 Numerical Experiments
We now test the performance of our proposed methods in a simulated stochastic shortest path problem following Hu et al. (2022), Elmachtoub and Grigas (2022). Specifically, we aim to go from the start node $s$ to the end node $t$ on a $5\times 5$ grid consisting of $d=40$ edges, where the costs of traveling on the edges are given by the random vector $Y\in\mathbb{R}^{40}$ (see Fig. 1). We consider covariates $X\in\mathbb{R}^{3}$ and a function $f_{0}(x)=\mathbb{E}\left[Y\mid X=x\right]$ whose components are cubic polynomials. The corresponding shortest path problem can be easily formulated into a CLO problem with the constraint set $\mathcal{Z}$ given by standard flow preservation constraints. The resulting optimal solution $z$ belongs to $\{0,1\}^{40}$, indicating whether passing each edge or not. We note that there are $m=70$ feasible paths $z_{1},\dots,z_{m}\in\mathcal{Z}$ from the source node to the target node, and the feasible paths are linearly dependent with a rank of $18$.
@@
We consider a bandit feedback setting, observing only the total traveling costs $C$ of the historical decisions $Z$ generated by a certain logging policy but not the edge-wise costs $Y$. We consider different types of logging policies that generate the decisions in the observed data. In this section, we report results using a random logging policy that picks a path from all feasible ones at random regardless of the covariate value. In Section 8, we further study the performance of two different covariate-dependent logging policies: one that picks paths according to the sign of the first covariate $X_{1}$, and one that depends on the signs of both $X_{1}$ and $X_{2}$. The empirical insights from the covariate-dependent logging policies are qualitatively the same as those for the random logging policy.
@@
We numerically evaluate the performance of the ETO approach and the IERM approach. For both approaches, we use the same class $\mathcal{F}$ to construct the decision policies. We consider three different classes for $\mathcal{F}$: a correctly specified polynomial class, a misspecified class that omits two high-order terms (termed degree-2 misspecification), and a misspecified class that omits four high-order terms (termed degree-4 misspecification). For the IERM approach, we consider DM and DR here but defer the results of ISW to Section 8.2 Table 3 for its significantly worse performance. In both DM and DR, the nuisance $f_{0}$ is estimated by a bandit-feedback regression given in Eq. 7 with a ridge penalty, and we test out the three aforementioned function classes for the nuisance class $\mathcal{F}^{\text{N}}$ as well. In DR, the nuisance $\Sigma_{0}(x)$ is estimated by the propensity score approach described in Remark 2.3, with the propensity scores estimated by either sample frequencies (for the random logging policy) or suitable decision tree models (for covariate-dependent logging policies). We further consider three variants when plugging $\hat{\Sigma}$ into the doubly robust score: pseudo-inverse (DR PI) as in the original $\theta_{\text{DR}}$ definition; lambda regularization (DR Lambda), where we replace $\Sigma^{\dagger}$ with $(\Sigma+\lambda I)^{-1}$ for a positive constant $\lambda$; and clipping (DR Clip), where eigenvalues of $\Sigma$ below a certain threshold are clipped to the threshold before taking the pseudo-inverse. For all IERM variants, we optimize the SPO+ losses as discussed in Section 4. Further details on the experimentation setup and implementation are summarized in Section 8. Finally, we consider some naive extensions of the offline contextual bandit learning with finite discrete actions (Naive ETO and Naive SPO+), where we view the feasible paths as separate discrete actions, without considering the linear structure of the decision-making problem. See Section 7 for details and Section 1.3 for background on offline bandits.
@@
We test the performance of different methods on an independent testing sample of size $2000$, and evaluate the ratio of their regrets relative to the expected cost of the global optimal policy $\pi_{f_{0}}$. Table 1 shows the average relative regret ratios of different methods across $50$ replications of the experiment for a random logging policy. Due to space limitations, we include results for the training data size $n=400,1000,1600$, and defer results for other sizes to Section 8.2. The relative regret of all methods properly decrease with the training data size $n$. In particular, the SPO+ approximation for the end-to-end IERM approach perform better than the ETO method. Among the SPO+ methods, the DM score achieves the best performance, while the DR score based pesudo-inverse performs the worst. Through a close inspection, we found that the bias adjustment term that involves the pseudo-inverse in the DR score causes a significant variance inflation. In fact, the ISW score also performs badly due to the high variance (see Section 8.2). The Lambda regularization and Clip techniques can effectively reduce the variance and result in improved decision-making. Moreover, the naive benchmarks that ignore the linear structure of the decision costs have much worse performance. This shows the importance of leveraging the linear problem structure.
@@
In Table 2, we show the performance of different methods when either the policy-inducing model $\mathcal{F}$ or the nuisance model $\mathcal{F}^{\text{N}}$ or both are misspecified. We observe that when the policy-inducing model is misspecified, the end-to-end SPO+ methods perform much better than the ETO method, provided that the nuisance model for SPO+ is correctly specified. This is consistent with the findings in the previous full-feedback CLO literature (e.g., Elmachtoub and Grigas 2022, Hu et al. 2022, Elmachtoub et al. 2023), showing the benefit of integrating estimation and optimization for misspecified policy models. However, the advantage of the end-to-end approaches is weakened dramatically once the nuisance model is misspecified. In this case, the evaluation of decision policies is biased, so the end-to-end approaches also target a “wrong” objective that may not accurately capture the decision quality. Moreover, we observe that when the nuisance model is misspecified, the DR score can somewhat outperform the DM score, because it can debias the misspecified nuisance to some extent. These results demonstrate new challenges with the bandit-feedback CLO: the end-to-end approaches are sensitive to the misspecification of nuisance models, and the DM and DR scores face different bias-and-variance tradeoffs under nuisance model misspecification. Therefore, in practice we may prefer more flexible models for accurate nuisance modeling, while using simpler policy-inducing models for tractable end-to-end optimization.

### v2

@@ 3.1 Main Theorem
Theorem 1.
Suppose Assumptions 2 and 3 hold, and $\mathcal{Z}^{*}(X)$ defined in Assumption 3 is a singleton almost surely. Suppose there exists a positive number $\tilde{r}$ (that depends on $n$) that upper bounds the critical radii of the function classes $\mathcal{G}_{1},\mathcal{G}_{2}$ almost surely (i.e., $\mathcal{R}_{n/2}(\mathcal{G}_{1},\tilde{r})\leq{\tilde{r}}^{2}$ and $\mathcal{R}_{n/2}(\mathcal{G}_{2},\tilde{r})\leq{\tilde{r}}^{2}$) and satisfies the inequalities $3n\tilde{r}^{2}/128\geq\log\log_{2}(1/\tilde{r})$, and $42\exp\left(-3n\tilde{r}^{2}/128\right)\leq\delta/$ 4. Then, there exists a positive constant $\tilde{C}(\alpha,\gamma)$ such that with probability at least $1-\delta$, we have
@@
The upper bound in Theorem 1 involves several different types of error terms. The first type is the critical radii $\tilde{r}$ that characterize the complexity of the function classes $\mathcal{G}_{1},\mathcal{G}_{2}$. This is a common complexity measure for function classes in statistics and machine learning [41]. For example, as we will discuss below, the critical radii scale as $\tilde{O}(\sqrt{\eta/n})$ if $\mathcal{G}_{1},\mathcal{G}_{2}$ are VC subgraph classes of dimension $\eta$. The second type is the term $\operatorname{Rate}^{\textbf{N}}(n/2,\delta/4)$ resulted from the errors in estimating the nuisance functions $f_{0},\Sigma_{0}$. Similar nuisance estimation errors also appear in the previous literature on offline contextual bandit learning [1, 45, 5] or more general learning problems that involve nuisance functions [13]. Below we will further discuss this term for different choices of score functions $\theta$. The third type is the misspecification error term $\text{Reg}\left(\tilde{\pi}^{*}\right)$. It is natural to expect a bigger decision regret when the misspecification error $\text{Reg}\left(\tilde{\pi}^{*}\right)$ is higher. In particular, when the policy class $\Pi_{\mathcal{F}}$ is well-specified, i.e., $\pi_{f_{0}}\in\Pi_{\mathcal{F}}$, we would have $\text{Reg}\left(\tilde{\pi}^{*}\right)=0$, and the regret upper bound would scale with the function class complexity through a fast rate $O(\tilde{r}^{\frac{2\alpha+2}{\alpha+2}})$. For VC subgraph classes with $\tilde{r}=\tilde{O}(\sqrt{\eta/n})$, the rate can range from $\tilde{O}(\sqrt{\eta/n})$ for $\alpha=0$ to $\tilde{O}({\eta/n})$ for $\alpha=\infty$. However, if the policy class is misspecified so that $\text{Reg}\left(\tilde{\pi}^{*}\right)>0$, then the dominating term related to $\tilde{r}$ would be a slow rate $O(\tilde{r})$. This reveals an interesting phase transition between the correctly specified and misspecified settings, which is not discovered in the previous theory that considers only well-specification [14].
Remark 3.
The constant coefficients in the regret upper bound can be improved when $2(1+\alpha)/\alpha$ is an integer (which accommodates the case of $\alpha=1$ justified in [15]):
@@
Notably, the constant in front of the misspecification error $\text{Reg}\left(\tilde{\pi}^{*}\right)$ becomes $1$ instead of $2$, which we believe is tight. This upper bound follows from nearly the same proof as Theorem 1 except that it handles an inequality slightly differently. Specifically, the proof of Theorem 1 involves a transcendental inequality of the form $\text{Reg}(\hat{\pi})\leq c_{1}\text{Reg}(\hat{\pi})^{\frac{\alpha}{2(1+\alpha)}}\tilde{r}+c_{2}$ for certain positive terms $c_{1},c_{2}$. This inequality is difficult to solve exactly, so we can only get an upper bound on its solution. It turns out that we can get a better upper bound when $2(1+\alpha)/\alpha$ is an integer.
@@
Proposition 2.
For any given $\delta\in(0,1)$, let $\chi_{n,\delta}$ be a positive sequence converging to 00 as $n\to\infty$, such that the mean square errors of the nuisance estimates satisfy the following with probability at least $1-\delta$:
@@
Compared to the DM and ISW scores, the impact of the estimation errors of the nuisances in the DR score is of only second order, i.e., $O(\chi_{n,\delta}^{2})$ instead of $O(\chi_{n,\delta})$. This echos the benefit of DR methods in causal inference and offline contextual bandit learning [1, 5, 6]. Notably, here we only need to bound the projected error on the nuisance estimator $\hat{f}$, which handles the setting when $\text{span}(\mathcal{Z})$ does not cover the whole $\mathbb{R}^{d}$ space, as is the case with the shortest path problem in Section 5.
@@ 3.2 Byproduct: Fast Rates in the Full Feedback Setting with Misspecification
Theorem 2.
Suppose $\mathcal{Z}^{*}(X)$ defined in Assumption 3 is a singleton almost surely. Define
@@
and $\tilde{r}^{\text{F}}$ be any solution to the inequality $\mathcal{R}_{n}(\mathcal{G}^{\text{F}},r)\leq r^{2}$ satisfying $3n(\tilde{r}^{\text{F}})^{2}/64\geq\log\log_{2}(1/\tilde{r}^{\text{F}})$ and $2\exp\left(-3n(\tilde{r}^{\text{F}})^{2}/64\right)\leq\delta$. If Assumption 3 further holds, then, with probability at least $1-\delta$,
@@
The regret bound is similar to that in Theorem 1 for the bandit setting, except that it does not have the nuisance error term $2\text{Rate}^{\textbf{N}}(n/2,\delta/4)$. This is because, in the full feedback setting, we observe the entire $Y$ vector, so we do not need to estimate any nuisance functions and can consider the nuisance estimation error term $\text{Rate}^{\textbf{N}}(n/2,\delta/4)$ as zero. When $\Pi_{\mathcal{F}}$ is a well-specified VC-subgraph class with dimension $\eta$, we have $\text{Reg}\left(\tilde{\pi}^{*}\right)=0$, and $\tilde{r}^{\text{F}}=\tilde{O}(\sqrt{\eta/n})$, so the bound in Theorem 2 reduces to $O((\eta/n)^{(\alpha+1)/(\alpha+2)}+\eta/n)$. This bound interpolates between $O(n^{-1/2})$ and $O(n^{-1})$ according to the margin parameter $\alpha$, recovering the fast rate in the full-feedback setting without misspecification as given in [14]. In contrast, our bound in Theorem 2 additionally quantifies the impact of policy misspecification, and its generalization Theorem 1 further incorporates the impact of nuisance estimation errors in the bandit-feedback setting.
@@ 4 Computationally Tractable Surrogate Loss
The IERM objective is generally nonconvex in $f\in\mathcal{F}$, making it computationally intractable to optimize. In the full feedback setting, tractable surrogate losses have been proposed [11, 16]. In this section, we briefly explain the SPO+ loss in [11] and how it can be used in the bandit setting.
@@
This IERM loss is equivalent to the “smart predict-then-optimize” (SPO) loss in Definition 1 of [11]. Letting $z^{*}(y)\in\arg\min_{z\in\mathcal{Z}}y^{\top}z$ with the same tie-breaking rule as in $\pi_{f}$, [11] propose the SPO+ surrogate loss:
@@ 5 Numerical Experiments
We now test the performance of our proposed methods in a simulated stochastic shortest path problem following [14, 11]. Specifically, we aim to go from the start node $s$ to the end node $t$ on a $5\times 5$ grid consisting of $d=40$ edges, where the costs of traveling on the edges are given by the random vector $Y\in\mathbb{R}^{40}$ (see Fig. 1). We consider covariates $X\in\mathbb{R}^{3}$ and a function $f_{0}(x)=\mathbb{E}\left[Y\mid X=x\right]$ whose components are cubic polynomials. The corresponding shortest path problem can be easily formulated into a CLO problem with the constraint set $\mathcal{Z}$ given by standard flow preservation constraints. The resulting optimal solution $z$ belongs to $\{0,1\}^{40}$, indicating whether passing each edge or not. We note that there are $m=70$ feasible paths $z_{1},\dots,z_{m}\in\mathcal{Z}$ from the source node to the target node, and the feasible paths are linearly dependent with a rank of $18$.
@@
We consider a bandit feedback setting, observing only the total traveling costs $C$ of the historical decisions $Z$ generated by a certain logging policy but not the edge-wise costs $Y$. We consider different types of logging policies that generate the decisions in the observed data. In this section, we report results using a random logging policy that picks a path from all feasible ones at random regardless of the covariate value. In Appendix C, we further study the performance of two different covariate-dependent logging policies: one that picks paths according to the sign of the first covariate $X_{1}$, and one that depends on the signs of both $X_{1}$ and $X_{2}$. The empirical insights from the covariate-dependent logging policies are qualitatively the same as those for the random logging policy.
@@
We numerically evaluate the performance of the ETO approach and the IERM approach [^1] The code scripts for all experiments can be found at https://github.com/CausalML/CLOBandit.. For both approaches, we use the same class $\mathcal{F}$ to construct the decision policies. We consider three different classes for $\mathcal{F}$: a correctly specified polynomial class, a misspecified class that omits two high-order terms (termed degree-2 misspecification), and a misspecified class that omits four high-order terms (termed degree-4 misspecification). For the IERM approach, we consider DM and DR here but defer the results of ISW to Section C.2 Table 3 for its significantly worse performance. In both DM and DR, the nuisance $f_{0}$ is estimated by a bandit-feedback regression given in Eq. 7 with a ridge penalty, and we test out the three aforementioned function classes for the nuisance class $\mathcal{F}^{\text{N}}$ as well. In DR, the nuisance $\Sigma_{0}(x)$ is estimated by the propensity score approach described in Remark 2, with the propensity scores estimated by either sample frequencies (for the random logging policy) or suitable decision tree models (for covariate-dependent logging policies). We further consider three variants when plugging $\hat{\Sigma}$ into the doubly robust score: pseudo-inverse (DR PI) as in the original $\theta_{\text{DR}}$ definition; lambda regularization (DR Lambda), where we replace $\Sigma^{\dagger}$ with $(\Sigma+\lambda I)^{-1}$ for a positive constant $\lambda$; and clipping (DR Clip), where eigenvalues of $\Sigma$ below a certain threshold are clipped to the threshold before taking the pseudo-inverse. For all IERM variants, we optimize the SPO+ losses as discussed in Section 4. Further details on the experimentation setup and implementation are summarized in Appendix C. Finally, we consider some naive extensions of the offline contextual bandit learning with finite discrete actions (Naive ETO and Naive SPO+), where we view the feasible paths as separate discrete actions, without considering the linear structure of the decision-making problem. See Appendix A for details and Section 1.3 for background on offline bandits.
@@
We test the performance of different methods on an independent testing sample of size $2000$, and evaluate the ratio of their regrets relative to the expected cost of the global optimal policy $\pi_{f_{0}}$. Table 1 shows the average relative regret ratios of different methods across $50$ replications of the experiment for a random logging policy. Due to space limitations, we include results for the training data size $n=400,1000,1600$, and defer results for other sizes to Section C.2. The relative regret of all methods properly decrease with the training data size $n$. In particular, the SPO+ approximation for the end-to-end IERM approach perform better than the ETO method. Among the SPO+ methods, the DM score achieves the best performance, while the DR score based pesudo-inverse performs the worst. Through a close inspection, we found that the bias adjustment term that involves the pseudo-inverse in the DR score causes a significant variance inflation. In fact, the ISW score also performs badly due to the high variance (see Section C.2). The Lambda regularization and Clip techniques can effectively reduce the variance and result in improved decision-making. Moreover, the naive benchmarks that ignore the linear structure of the decision costs have much worse performance. This shows the importance of leveraging the linear problem structure.
@@
In Table 2, we show the performance of different methods when either the policy-inducing model $\mathcal{F}$ or the nuisance model $\mathcal{F}^{\text{N}}$ or both are misspecified. We observe that when the policy-inducing model is misspecified, the end-to-end SPO+ methods perform much better than the ETO method, provided that the nuisance model for SPO+ is correctly specified. This is consistent with the findings in the previous full-feedback CLO literature [e.g., 11, 14, 12], showing the benefit of integrating estimation and optimization for misspecified policy models. However, the advantage of the end-to-end approaches is weakened dramatically once the nuisance model is misspecified. In this case, the evaluation of decision policies is biased, so the end-to-end approaches also target a “wrong” objective that may not accurately capture the decision quality. Moreover, we observe that when the nuisance model is misspecified, the DR score can somewhat outperform the DM score, because it can debias the misspecified nuisance to some extent. These results demonstrate new challenges with the bandit-feedback CLO: the end-to-end approaches are sensitive to the misspecification of nuisance models, and the DM and DR scores face different bias-and-variance tradeoffs under nuisance model misspecification. Therefore, in practice we may prefer more flexible models for accurate nuisance modeling, while using simpler policy-inducing models for tractable end-to-end optimization.